2 Math problems.. Having serious trouble. I really appreciate it!!!
1. Divide. Then simplify by taking roots, if possible. Assume all expressions under radicals represent positive numbers.
(Looks like a fraction)
On top 324xy under radical sign
On botton 3 (not to the third root, just a 3 before) then 3 under radical sign

2.Solve
x+8 = -3
The x+8 is under the radical sign..
Thank you!!!!!!


Answer by nyc_function(2741)

Post one question at a time to get a faster response.
sqrt{x + 8} = -3
We have a radical equation here.
(sqrt{x + 8})^2 = (-3)^2
x + 8 = 9
x = 9 - 8
x = 1
We must check to make sure that our answer for x satisfies the radical equation given to you.
In other words, it must have the same answer on both sides to satisfy the original radical equation.
So, let x = 1 and simplify.
sqrt{1 + 8} = -3
sqrt{9} = -3
3 = -3
Since we get different answers on both sides of the equation, the answer is NO SOLUTION.